# Learning from Reasoning Failures via Synthetic Data Generation

### v1

@@ 4.2 Synthetic data augmentation results
Table 2 provides in-domain evaluation results utilizing synthetic data derived from InfoVQA, ScienceQA, and OK-VQA. Notably, augmenting the LLaVA-Instruct dataset with our synthetic data achieves performance comparable to or better than using an equivalent amount of real domain-specific data in most cases. This result is particularly significant given that the synthetic samples were generated using only a small subset of the original training data: specifically, only those examples where LLaVA scored 0.0 while GPT scored 1.0. For instance, with OK-VQA, the original training set consists of 9,009 samples, but we utilized only 607 training samples which LLaVA failed on in order to generate 9009 synthetic samples, resulting in a performance boost of 13% on the OK-VQA test set. Similarly, our approach utilized only 28% of the ScienceQA training dataset to generate full synthetic replacements, yet still resulted in better performance than training directly on the real dataset. We also observe that performance improves as the amount of synthetic data used for data augmentation increases.
In practice, it may be desirable to combine synthetically generated data which was derived from reasoning failures across different datasets. We therefore provide results for two different sized mixtures of our synthetic data derived from InfoVQA, ScienceQA, and OK-VQA reasoning failures in Table 3. In addition to in-domain evaluations for these three datasets, we provide results for the five other datasets mentioned previously to measure the impact on OOD generalization. We also provide results for LLaVA models trained on three alternative synthetically generated datasets: ALLaVA, CoSyn-400k, and SimVQA [^1] ALLaVA consists of real images with synthetic text, whereas Cosyn-400k and SimVQA contain fully synthetic image-text examples. Our synthetic data utilized here contains both real and synthetic images.. From Table 3, we observe that our synthetic data outperforms almost all baselines (except VQAv2) in the maximum data augmentation setting ($N_{syn}=124,922$). When augmenting training sets with half as much data, we observe that our synthetic data produces the greatest improvements in-domain, whereas the ALLaVA dataset, which contains only real images paired with synthetic text, performs slightly better in other OOD settings. Finally, we note that training LLaVA in this mixed synthetic dataset setting produces similar in-domain performance as training LLaVA separately for reasoning failures derived from each individual dataset (Table 2).
Dataset Model $N$ $N_{syn}$ EM Score
InfoVQA LLaVA 624,610 0 26.7
@@
Table 2: In-domain evaluation of baseline LLaVA, LLaVA models trained using synthetically augmented data ($\textrm{LLaVA}_{\textrm{syn}}$) and using training data augmented with real in-domain data ($\textrm{LLaVA}_{\textrm{real}}$).
Base LLM $N$ $N_{syn}$ Augmentation Data TextVQA OCR-Bench InfoVQA OK-VQA VQAv2 ScienceQA MMBench MMMU
Vicuna-7B 624,610 0 N/A (baseline) 47.0 31.9 26.7 57.0 71.5 70.7 52.3 36.4
Vicuna-7B 687,071 62,461 ALLaVA 47.9 34.0 28.4 50.4 72.6 71.2 50.2 36.7
CoSyn-400K 47.1 31.8 28.5 55.8 72.3 71.5 52.4 34.6
SimVQA 46.8 31.6 26.6 54.4 71.8 71.1 53.5 34.7
Ours 47.4 33.2 33.1 60.8 71.3 73.1 52.5 36.2
Vicuna-7B 749,532 124,922 ALLaVA 47.2 34.1 28.8 49.4 72.8 66.5 43.5 34.2
CoSyn-400K 46.8 32.7 29.6 57.7 72.0 70.9 51.9 36.9
Ours 47.4 34.5 33.2 61.1 72.4 73.0 52.5 37.4
Gemma-2B 624,610 0 Ours 39.9 28.3 21.8 51.7 68.7 62.3 29.2 32.3
749,532 124,922 40.9 29.9 29.8 54.8 69.0 65.2 30.7 31.7
Qwen2-7B 624,610 0 Ours 45 31.4 26.7 59.2 71.7 77.9 63 42.3
749,532 124,922 46.2 32.5 27 60.6 72.9 80.7 63.5 42.2
@@
Table 3: Training data augmentation experimental results. $N$ denotes the total number of training examples, $N_{syn}$ denotes the number of synthetic examples in the training dataset generated using our approach. The first section of the table compared the same base LLM (Vicuna) trained on various datasets with our dataset, while the second section compares different LLM backbones trained on our dataset.
@@
To evaluate the generalization ability of our synthetic dataset, we trained models using the same datasets but with different backbone LLMs. In the following experiments, we used Gemma-2B [36] and Qwen2-7B [43] as base LLMs. We adopted the LLaVA two-phase training procedure: pretraining on the LLaVA 558k dataset [23] followed by instruction fine-tuning. Table 3 indicates that our synthetic data in its maximum augmentation setting, $N_{syn}=124,922$, outperform the baseline for both LLaVA-Gemma-2B and LLaVA-Qwen-7B almost on all benchmarks (except MMMU) Notably, despite being generated based on LLaVA-Vicuna-7B failure mechanisms, our synthetic data enhance the performance of other models whether they are of same size (Qwen2-7B) or smaller (Gemma-2B).
@@ 4.3 Correcting specific types of reasoning failures
Figure 4: Figure shows the clusters of LLAVA reasoning failures described by GPT-4o.
Our synthetic data generation approach explicitly identifies different types of LMM reasoning failures. To systematically categorize these failures, we encoded each reasoning explanation using sentence transformers  [33] and clustered them using k-means. Figure  4 presents the resulting clusters, highlighting prevalent failure modes such as optical character recognition (OCR) and object detection errors. Based on this analysis, we further investigated whether targeted synthetic data can effectively address these specific failure cases and enhance LLAVA’s reasoning capabilities.
Specifically, we augmented LLaVA-Instruct with 10,579 synthetic samples from our $\textrm{VizWiz}_{\textrm{syn}}$-MFS addressing object detection reasoning failures and repeated the second stage of LLaVA finetuning. The model was then evaluated on CIFAR-10  [17], Food-101  [2], iNaturalist  [38], MNIST  [19], Fashion-MNIST  [41], and Oxford-Pets (Binary)  [32] by formatting samples as multiple-choice questions. Table  1 presents a comparison of LLaVA and $\textrm{LLaVA}_{\textrm{syn}}$. The results show that $\textrm{LLaVA}_{\textrm{syn}}$ surpasses LLaVA on four out of six datasets, with particularly notable improvements on iNaturalist, MNIST and Oxford-Pets. This demonstrates the significant impact of our synthetic dataset in addressing specific reasoning failures within LLAVA. By systematically incorporating targeted synthetic samples, we can mitigate common failure cases, leading to measurable performance improvements across multiple benchmarks. Our findings highlight the effectiveness of leveraging targeted synthetic data to refine model reasoning and suggest that incorporating such data-driven interventions can significantly enhance the robustness and generalization of LMMs.
@@ 5.1 Human evaluation of dataset quality
Three of the authors of this work conducted a human evaluation by assessing three different aspects of our generated samples: (1) the alignment of the question and answer in relation to the image prompt, (2) the alignment between the image prompt and the generated image, and (3) the correctness of the answer given the question and image. The first evaluation reflects the quality of reasoning, the second evaluates the fidelity of the image generator’s output, and the third combines both aspects. Scores range from 1 to 3, where 1 indicates an irrelevant alignment, 3 signifies a relevant alignment, and 2 represents a partially relevant or ambiguous alignment. We evaluated 200 samples in total, with 101 containing real images and 99 being fully synthetic. The overall correctness score for answers was 2.78, with real‐image‐based samples scoring 2.75 and fully synthetic samples scoring 2.81, indicating that fully synthetic samples achieve a level of fidelity equal to or even slightly exceeding that of real‐image-based samples. For the synthetic samples specifically, we also measured the alignment between the image prompt and the generated image (2.66), and the alignment of the generated question and answer with the image prompt (2.84), indicating the high quality of reasoning in the generated responses.
5.2 Training data substitution vs. augmentation and impact of synthetically generated images
Train Data $N$ $N_{syn}$ TextVQA OCR-Bench InfoVQA VQAv2 OK-VQA ScienceQA MMBench MMMU
Baseline 624,610 0 47.0 31.9 26.7 71.5 57.0 70.7 52.3 36.4
624,610 156153 46.9 31.1 27.0 72.7 57.0 70.6 51.2 37.9
687,071 62,461 47.7 31.8 25.8 71.6 59.4 71.2 52.3 36.4
687,071 62,461 47.2 32.2 27.2 71.8 56.9 71.2 52.3 33.8
Table 4: Ablation experiments comparing baseline LLaVA to LLaVA models trained with synthetic data generated from VizSiz failures. We investigate substitution and augmentation strategies for synthetic data, as well as the use of synthetic vs. real images.
Our previous experiments augmented an existing 624k sample training dataset (LLaVA-Instruct) with our synthetic data. In domains where data is scarce, training datasets of this size may not be available. To investigate the utility of our synthetic data in such low-resource settings, we conducted experiments in which we randomly substituted different quantities of examples from the original dataset with our synthetically generated data [^2] We used synthetic data derived from Vizwiz failures in this setting.. The results of this experiment are provided in rows 2-3 of Table 4. Even when up to 25% of the original dataset is substituted with our synthetic data, we achieve performance that is either as good or better than the baseline LLaVA model across a broad range of downstream tasks. This is despite the fact that the original LLaVA training dataset utilizes real images, whereas our synthetic data used in this experiment contained only synthetically generated images. The fact that our synthetic data achieves similar or better performance than an existing real data source is significant, as prior studies have shown that training on synthetically generated image data is often much less efficient than training on an equivalent amount of real image data [12]. Table 4 also shows the impact of using real vs. synthetic images in our pipeline. Specifically, we compare the effectiveness of our synthetic data derived from Vizwiz reasoning failures when paired with real images (from Vizwiz) or synthetically generated images. In the training data augmentation setting, we observe that synthetic images generally achieve similar results as utilizing real images. Synthetic images even surpass the performance of real images in TextVQA, OK-VQA, and MMBench. This demonstrates the high quality of our synthetic images and their potential to serve as replacements for real images in low-resource settings where data is scarce.
5.3 Impact of filtering on data quality
5.4 Comparison of LLM synthetic data generators
We compared two frontier LMMs, GPT-4o and Qwen2-VL-7B [40], for generating synthetic data grounded in LLaVA-7B failures. Qwen2-VL-7B was chosen for its high accuracy on vision-language benchmarks. However, samples generated by Qwen2-VL reduced downstream performance by 2% on InfoVQA and 6.5% on OK-VQA compared to GPT-4o. Additionally, those samples received lower filtering scores: 1.9 (Qwen2-VL) vs. 2.5 (GPT-4o) for OK-VQA, and 2.6 vs. 2.9 for InfoVQA. Based on our manual analysis, we hypothesiize that these differences may result from the detailed and precise reasoning provided by GPT-4o, resulting in synthetic samples that are better tailored to address identified reasoning failures. In contrast, samples generated by Qwen2-VL-7B sometimes demonstrate lower diversity, which could limit their effectiveness in addressing the broad range of failure modes. Figure 7 (Section 9.1 of the supplementary material) provides an example of these observed differences in the reasoning processes of GPT-4o and Qwen2-VL-7B models, as well as the corresponding generated fully synthetic samples.
@@ 6 Conclusion
We developed a novel method for generating multimodal synthetic data by analyzing model reasoning failures. This approach enabled us to create a multimodal instruction tuning dataset with over 553k synthetic examples derived from LLaVA’s errors. Experimental results show that our synthetic data significantly improves LLaVA’s in-domain performance on InfoVQA, ScienceQA, and OK-VQA, even outperforming training on an equivalent amount of real data sampled from most of these datasets. Furthermore, models trained on our synthetic dataset exhibit improvements in OOD evaluations and outperform training on other existing synthetic datasets when the amount of training data augmentation is scaled. We also showed that training LLaVA only on examples derived from specific failure modes improves its performance on tasks which require corresponding forms of reasoning. Through additional ablations and human evaluations, we demonstrated the effectiveness of different components of our methodology and the high quality of our dataset. We believe our study points to the promise of targeted synthetic data generation strategies which leverage an understanding of a model’s reasoning deficiencies to construct better synthetic examples.
Supplementary Material
7 Implementation details
7.1 Compute Infrastructure

### v2

@@ 4.2 Synthetic Data Augmentation Results
Table 1 provides in-domain evaluation results utilizing synthetic data derived from InfoVQA, ScienceQA, and OK-VQA. Notably, augmenting the LLaVA-Instruct dataset with our synthetic data achieves performance comparable to or better than using an equivalent amount of real domain-specific data in most cases. This result is particularly significant given that the synthetic samples were generated using only a small subset of the original training data: specifically, only those examples where LLaVA scored 0.0 while GPT scored 1.0. For instance, in OK-VQA, the original training set consists of 9,009 samples, but we utilized only 607 training samples where LLaVA failed in order to generate 9009 synthetic samples, resulting in a performance boost of 13% on the OK-VQA test set. Similarly, our approach utilized only 28% of the ScienceQA training dataset to generate full synthetic replacements, yet still resulted in better performance than training directly on the real dataset. We also observe that performance improves as the amount of synthetic data used for data augmentation increases.
In practice, it may be desirable to combine synthetically generated data which was derived from reasoning failures across different datasets. We therefore provide results for two different sized mixtures of our synthetic data derived from InfoVQA, ScienceQA, and OK-VQA reasoning failures in Table 2. We also provide results for LLaVA models trained on four alternative synthetically generated datasets: ALLaVA (Chen et al. 2024b), CoSyn-400k (Yang et al. 2025), SimVQA (Cascante-Bonilla et al. 2022), and Img-Diff (Jiao et al. 2024) [^1] ALLaVA consists of real images with synthetic text, whereas Cosyn-400k and SimVQA contain fully synthetic examples. Our synthetic data utilized here contains both real and synthetic images.. From Table 2, we observe that our synthetic data outperforms these baselines datasets, demonstrating the value of grounding synthetic data generation in an analysis of a model’s failures. Finally, we evaluated trained models on OOD benchmarks (Table 8 of Appendix) and found that models trained on our dataset do not exhibit degradation in OOD performance, demonstrating the ability of our approach to correct model reasoning failures without sacrificing general reasoning capabilities. Importantly, our dataset strikes a middle ground in the cost–quality–diversity tradeoff and can be further complemented with lower-cost image generation pipelines such as CoSyn’s to achieve scalable and semantically rich supervision. A detailed comparison of compute and generation costs across our method and baseline datasets is provided in Appendix B.3.
Dataset Model NN $N_{syn}$ EM Score
I-VQA LLaVA 624,610 0 26.7
@@
Table 1: In-domain evaluation of baseline LLaVA, LLaVA trained using synthetically augmented data ($\textrm{LLaVA}_{\textrm{syn}}$), and augmented with real in-domain data ($\textrm{LLaVA}_{\textrm{real}}$). I-VQA denotes InfoVQA and SQA denotes ScienceQA.
Base LLM NN $N_{syn}$ Aug Data I-VQA OK-VQA SQA
Vicuna-7B 625k 0 N/A (bs) 26.7 57.0 70.7
Vicuna-7B 643k 18k Img-Diff 27.3 56 71.3
Ours 31.5 61.2 72
Vicuna-7B 687k 62k SimVQA 26.6 54.4 71.1
Ours 33.1 60.8 73.1
Vicuna-7B 750k 125k ALLaVA 28.8 49.4 66.5
CoSyn 29.6 57.7 70.9
Ours 33.2 61.1 73.0
Gemma-2B 625k 0 Ours 21.8 51.7 62.3
749k 125k 29.8 54.8 65.2
Qwen2-7B 625k 0 Ours 26.7 59.2 77.9
750k 125k 27 60.6 80.7
@@
Table 2: Training data augmentation results. NN denotes the total number of training examples, $N_{syn}$ denotes the number of synthetic examples in the training dataset generated using our approach. I-VQA denotes InfoVQA and SQA represents ScienceQA. The first section of the table compared the same base LLM (Vicuna) trained on various datasets with our dataset, while the second section compares different LLM backbones trained on our dataset.
4.3 Generalization to Other Models and Tasks
@@
While our synthetic data is collected by targeting the reasoning failures of a specific LMM, we also evaluated whether it could generalize to other models and tasks, thereby further increasing its value beyond correcting targeted reasoning failures in a single model. To evaluate this generalization ability of our synthetic dataset, we trained models using the same datasets as before but with different backbone LLMs. In the following experiments, we used Gemma-2B (Team et al. 2024) and Qwen2-7B (Yang et al. 2024) as base LLMs. We adopted the LLaVA two-phase training procedure: pretraining on the LLaVA 558k dataset (Liu et al. 2024a) followed by instruction fine-tuning. Table 2 indicates that our synthetic data in its maximum augmentation setting, $N_{syn}=124,922$, outperform the baseline for both LLaVA-Gemma-2B and LLaVA-Qwen-7B across nearly all benchmarks. Notably, despite being generated based on LLaVA-Vicuna-7B failure modes, our synthetic data enhances the performance of other models whether they are of the same size (Qwen2-7B) or smaller (Gemma-2B).
For some models (e.g., Qwen2.5-VL-7B-Instruct), the original finetuning dataset is not publicly available. Nevertheless, we show that our dataset can be sufficient to improve these models through continued finetuning. Table 4 provides the results of finetuning Qwen2.5-VL-7B on our 124K synthetic-only dataset derived from LLaVA reasoning failures. We observe that finetuning Qwen2.5-VL-7B on our 124K dataset improves performance significantly on OK-VQA and ScienceQA. InfoVQA performance remains unchanged, likely due to Qwen’s high baseline and prior exposure to similar data. Generating synthetic data which targets Qwen-specific failure cases may yield further gains; nevertheless, these results point to the value of our dataset beyond correcting reasoning failures in only the targeted LMM.
Finally, we investigated whether our targeted synthetic dataset could improve performance on other tasks besides those used for deriving the reasoning failures. Specifically, we evaluated our $\textrm{LLaVA}_{\textrm{syn}}$ model trained on the 124K synthetic dataset on multiple hallucination and math benchmarks. The results in Table 3 show that $\textrm{LLaVA}_{\textrm{syn}}$ outperforms the baseline LLavA model across all such benchmarks, demonstrating how our synthetic dataset improves model performance beyond the specific reasoning failures targeted during the dataset generation process. The ability of our synthetic dataset to generalize to improving other LMMs and tasks than the ones used in the generation process further illustrates the significant value of our approach.
LLaVA $\textrm{LLaVA}_{\textrm{syn}}$ (124k)
[…]
LLaVA-B-$CO\uparrow$ 85.3 86.9
[…]
Table 3: Our targeted 124k synthetic dataset generalizes to improve LLaVA on hallucination and math benchmarks.
4.4 Continued Finetuning Results
We also conducted experiments wherein we continued finetuning LLaVA on our 124K mixture dataset. This approach allows us to evaluate performance gains attributed solely to our data while significantly reducing compute requirements compared to full finetuning. In this setting, we used the same hyperparameter configurations as our previous finetuning experiments. Table 4 shows that continued finetuning improves in-domain performance except on OK-VQA. Notably, continued finetuning on the real OK-VQA training set (9K, 41.1%) results in worse performance than with an identical-size synthetic dataset (48.2%), suggesting possible overfitting when training on real data.
InfoVQA OK-VQA SciQA
LLaVA1.5-7b 26.4 57.1 70.8
+ CFT 32.4 45.8 73.2
Qwen2.5-VL-7b-Instruct 82.6 42.4 67.9
+ CFT 82.5 53.0 85.5
Table 4: Continued finetuning results for LLaVA-1.5-7b and Qwen2.5-VL-7B-Instruct with our 124K synthetic dataset
@@ 4.5 Correcting Specific Types of Reasoning Failures
Cluster name Percentage (%)
Feature misinterpretation 8
Blurriness 12.4
Weakness in visual analysis 18.6
Incomplete context understanding 24
Other 11.3
Text recognition errors 9.2
Object recognition failure 10.8
Overgeneralization 5.6
Table 5: Table shows the clusters of LLAVA reasoning failures described by GPT-4o.
Our synthetic data generation approach identifies specific types of LMM failures. To systematically categorize these failures, we encoded each reasoning explanation using sentence transformers  (Reimers and Gurevych 2019) and clustered them using k-means. Table  5 presents the resulting clusters, highlighting prevalent failure modes such as optical character recognition (OCR) and object detection errors. Based on this analysis, we further investigated whether targeted synthetic data can effectively address these specific failure cases and enhance LLAVA’s reasoning capabilities.
Dataset LLAVA $\textrm{LLaVA}_{\textrm{syn}}$
CIFAR-10 82.1 81.2
Food-101 13.4 13.2
iNaturalist 20.6 52.0
MNIST 75.1 80.5
F-MNIST 9.8 10.0
Oxford-pets 39.6 96.4
Table 6: Image classification results, with $\textrm{LLaVA}_{\textrm{syn}}$ augmented only using examples for object recognition failures.
Specifically, we augmented LLaVA-Instruct with 10,579 synthetic samples from our $\textrm{VizWiz}_{\textrm{syn}}$-MFS addressing object detection reasoning failures and repeated the second stage of LLaVA finetuning. The model was then evaluated on CIFAR-10 (Krizhevsky et al. 2009), Food-101 (Bossard et al. 2014), iNaturalist (Van Horn et al. 2018), MNIST (LeCun et al. 1998), Fashion-MNIST (Xiao et al. 2017), and Oxford-Pets (Binary) (Parkhi et al. 2012) by formatting samples as multiple-choice questions. Table 6 presents a comparison of LLaVA and $\textrm{LLaVA}_{\textrm{syn}}$. $\textrm{LLaVA}_{\textrm{syn}}$ surpasses LLaVA on four out of six datasets, with significant improvements on iNaturalist, MNIST and Oxford-Pets. This demonstrates the effectiveness of our synthetic dataset in addressing specific reasoning failures within LLaVA. Our findings highlight the effectiveness of leveraging targeted synthetic data to refine model reasoning and suggest that incorporating such data-driven interventions can significantly enhance the robustness and generalization of LMMs.
@@ 5.1 Human Evaluation of Dataset Quality
Three of the authors of this work conducted a human evaluation by assessing three different aspects of our generated samples: (1) the alignment of the question and answer in relation to the image prompt, (2) the alignment between the image prompt and the generated image, and (3) the correctness of the answer given the question and image. The first evaluation reflects the semantic coherence, the second evaluates the fidelity of the image generator’s output, and the third combines both aspects. Scores range from 1 to 3, where 1 indicates an irrelevant alignment, 3 signifies a relevant alignment, and 2 represents a partially relevant or ambiguous alignment. We evaluated 200 samples in total, with 101 containing real images and 99 being fully synthetic. The overall correctness score for answers was 2.78, with real‐image‐based samples scoring 2.75 and fully synthetic samples scoring 2.81, indicating that fully synthetic samples achieve a level of fidelity equal to or even slightly exceeding that of real‐image-based samples. For the synthetic samples specifically, we also measured the alignment between the image prompt and the generated image (2.66), and the alignment of the generated question and answer with the image prompt (2.84), indicating the high quality of reasoning in the generated responses. Additionally, to assess the frontier model’s reasoning and ensure it highlights meaningful failures, five annotators rated 209 samples (image, QA, and error analysis) on a 1–3 scale. The average score of 2.81 indicates that the explanations usually identified the true failure cause.
5.2 Impact of LMM on Generated Data
We evaluate our synthetic‐data pipeline with two LMMs of very different cost profiles: the proprietary GPT-4o API and the fully open, Apache-2.0-licensed Qwen2-VL-7B (Wang et al. 2024). Both models improve LLaVA-7B after fine-tuning on the resulting data: GPT-4o yields gains of 2.7% on InfoVQA and 10% on OK-VQA, while Qwen2-VL-7B delivers 7.0% and 3.5% on the same tasks. Although GPT-4o’s detailed and precise reasoning may contribute to generating more targeted and effective synthetic samples, which produces larger gains on some benchmarks, our failure-guided generation shows robustness by producing meaningful improvements across both models. Because Qwen2-VL-7B runs locally on a single A100 GPU, synthetic data can be obtained with our approach at a lower cost than commercial models such as GPT-4o. Figure 7 (Section C.1 in Appendix) illustrates representative examples of the reasoning and sample diversity across the two models.
@@ 6 Conclusion
We introduced a novel method for generating multimodal synthetic data by analyzing model reasoning failures. This approach enabled us to create a multimodal instruction tuning dataset with over 553k synthetic examples derived from LLaVA’s errors. Experimental results show that our synthetic data significantly improves LLaVA’s in-domain performance on InfoVQA, ScienceQA, and OK-VQA, even outperforming training on an equivalent amount of real data sampled from most of these datasets. Furthermore, models trained on our synthetic dataset exhibit improvements in OOD evaluations and outperform training on other existing synthetic datasets when the amount of training data augmentation is scaled. We also showed that training LLaVA only on examples derived from specific failure modes improves its performance on tasks which require corresponding forms of reasoning. Through additional ablations and human evaluations, we demonstrated the effectiveness of different components of our methodology and the high quality of our dataset. We believe our study points to the promise of targeted synthetic data generation strategies which leverage an understanding of a model’s reasoning deficiencies to construct better synthetic examples.
Appendix A Implementation details
A.1 Compute Infrastructure